### v1

# Polarization-Dependent Loss Mitigation with Orthogonal-Design-Based Precoding and Interference Cancellation

Abstract
Recent work by Shehadeh and Kschischang provides a simple capacity-achieving scheme for channels with polarization-dependent loss (PDL) under common modeling assumptions via a careful choice of orthogonal-design-based precoding and interference cancellation. This letter extends that work with a simulation-based demonstration showing that this scheme remains highly effective at mitigating PDL in the highly practical setting of 4-PAM with Chase-decoded extended Hamming inner codes rather than the near-capacity inner codes considered in the original work. An alternative near-optimal variation of this scheme is provided requiring only one inner code rather than two and suffering no penalty in the absence of PDL, making it much more practical.
Index Terms:
Optical fiber communication, successive interference cancellation, polarization-division multiplexing, polarization-dependent loss.
publicationid: pubid: $0$–$00$ 2025 IEEE
I Setup
We consider a memoryless model for a polarization-division-multiplexed (PDM) coherent optical communication system with slowly-varying polarization-dependent loss (PDL) and deterministic insertion loss. We further assume that the channel parameters are perfectly known to the receiver but unknown to the transmitter. Such a model is considered in [1, 2, 3, 4, 5, 6, 7, 8, 9] and it is shown by Shehadeh and Kschischang in [1] that a certain orthogonal-design-based precoding and interference cancellation scheme is capacity-achieving under this model.
We further assume in this letter real-valued channels with real-valued inputs. However, we emphasize that everything presented generalizes to complex-valued channels and inputs by analogously considering the complex variant of the scheme included in [1]. We then have a two-parameter channel class
$\underbrace{\begin{bmatrix}Y_{1}\\ Y_{2}\end{bmatrix}}_{\mathbf{Y}}=\underbrace{\begin{bmatrix}\sqrt{1+\gamma}&0\\ 0&\sqrt{1-\gamma}\end{bmatrix}}_{\mathbf{D}_{\gamma}}\underbrace{\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}}_{\mathbf{R}_{\theta}}\underbrace{\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}}_{\mathbf{X}}+\underbrace{\begin{bmatrix}Z_{1}\\ Z_{2}\end{bmatrix}}_{\mathbf{Z}}$
where $\gamma\in[-\alpha,\alpha]$ and $\theta\in[0,2\pi)$, representing an adversarial class of channels with two polarizations and up to
$10\log_{10}\left(\frac{1+\alpha}{1-\alpha}\right)\text{ dB}$
of PDL where $\alpha\in[0,1)$ is a fixed parameter. Moreover, $\mathbf{X}$ and $\mathbf{Z}$ are independent with $\mathbf{Z}$ being standard white Gaussian and $\mathbf{X}$ satisfying $\mathbb{E}[\norm{\mathbf{X}}_{2}^{2}]=2\cdot\mathsf{SNR}$ where $\mathsf{SNR}$ denotes the signal-to-noise ratio (SNR). We are generally interested in the worst-case performance of any given scheme over all $\gamma\in[-\alpha,\alpha]$ and $\theta\in[0,2\pi)$.
When precoding, we consider a two-channel-use extension of the channel matrix given by $\operatorname{diag}(\mathbf{D}_{\gamma}\mathbf{R}_{\theta},\mathbf{D}_{\gamma}\mathbf{R}_{\theta})\in\mathbb{R}^{4\times 4}$ which we can interpret either as representing two uses of the channel in time or the use of a complex-valued input with real-valued representations where the first and second halves of a vector represent the real and imaginary parts respectively. We then take the input to the channel to be $\mathbf{X}=\mathbf{G}\mathbf{U}$ where $\mathbf{X},\mathbf{U}\in\mathbb{R}^{4\times 1}$ and $\mathbf{G}\in\mathbb{R}^{4\times 4}$ is an orthogonal precoding matrix. The effective channel is then described by the effective channel matrix $\operatorname{diag}(\mathbf{D}_{\gamma}\mathbf{R}_{\theta},\mathbf{D}_{\gamma}\mathbf{R}_{\theta})\mathbf{G}$ and has input $\mathbf{U}$ satisfying $\mathbb{E}[\norm{\mathbf{U}}_{2}^{2}]=2\cdot\mathsf{SNR}$.
We use bit-interleaved coded modulation (BICM) [10] with extended Hamming codes and a Gray-labeled $4$-PAM constellation together with soft-decision Chase decoding [11]. Since the scheme of [1] requires two codes with different rates used in equal proportion, we choose them to have the same average rate as a third code which is used in schemes requiring one code. In particular, we consider three codes with lengths and dimensions $(n_{1},k_{1})$, $(n_{2},k_{2})$, and $(n,k)$ respectively satisfying
$\frac{1}{2}\left(\frac{k_{1}}{n_{1}}+\frac{k_{2}}{n_{2}}\right)=\frac{k}{n}\text{.}$
We take $(n_{1},k_{1})=(100,90)$, $(n_{2},k_{2})=(400,388)$, and $(n,k)=(200,187)$. These codes are obtained by shortening longer extended Hamming parent codes. This yields a rate of $0.935$ so that the inner coding scheme is somewhat comparable to that in the 400ZR implementation agreement [12].
Our key contributions are the demonstration of the effectiveness of the scheme of [1] for PDL mitigation in this context and the introduction of a variation on this scheme with various practical advantages. We illustrate our points by comparing five schemes defined in Section II which include the scheme of [1] and the proposed variation. This is followed by simulations in Section III and concluding remarks in Section IV.
II Five Schemes
We now describe five schemes which we identify as iZ, pZ, D, pD, and the scheme of [1]. These are based on precoding, linear minimum mean square error (LMMSE) equalization, and successive interference cancellation (SIC) [13, Chapter 8]. We use zero-forcing (ZF) instead of LMMSE for simplicity since it is comparable at the high SNRs of interest. Implementation details for all schemes can be found in our complete simulation code which we make available online [14].
II-A One Code, Spatio-Temporal Interleaving, and ZF without Precoding (iZ)
Scheme iZ is our baseline scheme in which we interleave two codewords from one code across the two polarizations and in time to decorrelate the noise. We then equalize via ZF and decode each codeword independently. This is easy to do and achieves some PDL penalty mitigation relative to treating each polarization completely independently.
II-B One Code and ZF with Precoding (pZ)
In Scheme pZ, we again send two codewords from one code but consider a two-channel-use extension and the $4\times 4$ precoding matrix from [1]:
$\mathbf{G}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&0&1&0\\ 0&1&0&1\\ 0&1&0&-1\\ -1&0&1&0\end{pmatrix}\text{.}$ (1)
This is followed by ZF and independent decoding of each codeword which sees a worst-case effective gain of $1-\alpha^{2}$. The resulting scheme is comparable to those of [15, 9].
II-C One Code and D-BLAST-Style ZF-SIC without Precoding (D)
In Scheme D, we consider a D-BLAST-style [13, Chapter 8] approach which allows us to perform ZF-SIC on one codeword by diagonally staggering it in time and space. After mapping to a $4$-PAM constellation, we have a codeword of $n/2$ symbols which is split into two halves $\mathbf{x}^{\mathsf{f}}$ and $\mathbf{x}^{\mathsf{s}}$ of $n/4$ symbols. We then transmit $\zeta$ such codewords in $(\zeta+1)n/4$ channel uses in time which we represent by the $42\times(\zeta+1)n/$ 4 matrix
$\begin{pmatrix}\psi&\mathbf{x}_{1}^{\mathsf{s}}&\mathbf{x}_{2}^{\mathsf{s}}&\cdots&\mathbf{x}_{\zeta-2}^{\mathsf{s}}&\mathbf{x}_{\zeta-1}^{\mathsf{s}}&\mathbf{x}_{\zeta}^{\mathsf{s}}\\ \mathbf{x}_{1}^{\mathsf{f}}&\mathbf{x}_{2}^{\mathsf{f}}&\mathbf{x}_{3}^{\mathsf{f}}&\cdots&\mathbf{x}_{\zeta-1}^{\mathsf{f}}&\mathbf{x}_{\zeta}^{\mathsf{f}}&\psi\end{pmatrix}$
where $\psi$ denotes some fixed symbol sequence known at the receiver and transmitter. This results in an overall rate of
$\frac{\zeta}{\zeta+1}\cdot\frac{k}{n}$
with $\zeta$ controlling the rate loss at the cost of increased impact of error propagation. We then decode $\zeta$ codewords by repeating the following steps for each column in the above:
  • If this is the first column, cancel the interference corresponding to $\psi$; otherwise, cancel the interference corresponding to the second half estimate obtained in the previous step. Equalize after interference cancellation to produce a noisy first half for the next step.
  • Equalize the next column, append the second half of our noisy codeword to the noisy first half obtained in the previous step, and decode it. We use the second half of the decoded codeword as the estimate for the next step.
What this accomplishes relative to iZ is that the first half of every codeword always sees a cleaner interference-free channel. However, this usually cannot improve upon iZ in this setting since the case of $\theta=0$, which is already interference-free, is usually the worst case. In particular, for both iZ and D, when $\theta=0$ and $\gamma=\pm\alpha$, the two halves of each codeword see effective gains of $1+\alpha$ and $1-\alpha$. This is usually the worst case for off-the-shelf coded modulation schemes which are not designed for gain imbalance. This leads to the next scheme.
Figure 1: Simulation results for Schemes iZ, pZ, D, pD and the scheme of [1] for $\alpha=0$ and $\alpha=0.599$ corresponding respectively to worst-case PDLs of 00 dB and $6$ dB.
II-D One Code and D-Blast-Style ZF-SIC with Precoding (pD)
In Scheme pD, we consider combining the orthogonal-design-based precoding and interference cancellation approach of [1] with a D-BLAST-style technique as in Scheme D so that only a single code is used. In particular, we consider a two-channel-use extension and the precoding matrix (1) after which we apply Scheme D to the resulting effective channel. After mapping to a $4$-PAM constellation, we have a codeword of $n/2$ symbols which we split into two halves $\mathbf{u}^{\mathsf{f}}$ and $\mathbf{u}^{\mathsf{s}}$ of dimensions $82\times n/$ 8 containing $n/4$ symbols each. We then transmit $\zeta$ such codewords in $(\zeta+1)n/8$ uses of the two-channel-use extended channel so that the input is represented by the $84\times(\zeta+1)n/$ 8 matrix
$\mathbf{G}\begin{pmatrix}\psi&\mathbf{u}_{1}^{\mathsf{s}}&\mathbf{u}_{2}^{\mathsf{s}}&\cdots&\mathbf{u}_{\zeta-2}^{\mathsf{s}}&\mathbf{u}_{\zeta-1}^{\mathsf{s}}&\mathbf{u}_{\zeta}^{\mathsf{s}}\\ \mathbf{u}_{1}^{\mathsf{f}}&\mathbf{u}_{2}^{\mathsf{f}}&\mathbf{u}_{3}^{\mathsf{f}}&\cdots&\mathbf{u}_{\zeta-1}^{\mathsf{f}}&\mathbf{u}_{\zeta}^{\mathsf{f}}&\psi\end{pmatrix}$
where $\psi$ is some fixed $82\times n/$ 8 symbol sequence. As before, we have an overall rate of $\zeta/(\zeta+1)\cdot k/n$. We then decode exactly as in Scheme D but we consider the effective channel combined with $\mathbf{G}$ for the purposes of equalization and interference cancellation.
Unlike the case of Scheme D, the two halves of each codeword respectively see effective gains of $1$ and $1-\alpha^{2}$. This is due to the fact that (1) induces an orthogonal design [16, 17] in each half of the effective channel matrix as shown in [1]. While we expect a gain over pZ since half of our symbols have a strictly improved reliability, this scheme will still be usually sub-optimal due to the remaining (but reduced) gain imbalance.
II-E Two Codes and ZF-SIC with Precoding ([1])
Lastly, we consider the scheme proposed in [1] using two codes with average rate $k/n$ as described in Section I. This entails precoding as in pZ but sending two codewords from the two different codes and using ZF-SIC rather than ZF. Due to the orthogonal design associated with (1), the lower rate and thus stronger code always sees a worst-case gain of $1-\alpha^{2}$ while the higher rate and thus weaker code always sees a worst-case gain of $1$. As shown in [1], the resulting scheme is strictly optimal in the sense that the performance is only limited by the classical additive white Gaussian noise channel performance of the two codes used and the fundamental information-theoretic cost of PDL.
III Simulations
We now provide bit error rate (BER) versus SNR simulation results for each of the five schemes. For Schemes D and pD, we take $\zeta=100$. Chase decoding parameters and other details can be found in our simulation code [14]. We plot in Fig. 1 the worst-case (maximum) BER over all $\gamma\in[-\alpha,\alpha]$ and $\theta\in[0,2\pi)$ (which we discretize) for $\alpha=0$ and $\alpha=0.599$ corresponding respectively to channel classes with a worst-case PDL of 00 dB and a worst-case PDL of $6$ dB.
Fig. 1 shows that pD is worse than the scheme of [1] by about $0.25$ dB at a BER of $10^{-2}$ and by about $0.125$ dB at a BER of $10^{-3}$. However, the scheme of [1] alone suffers a significant penalty in the absence of PDL ($\alpha=0$) for BERs below $10^{-2}$ relative to all other schemes considered. This is because the two different rates of the two codes used must be tuned according to the worst-case PDL value. In practice, we may wish for a single system to perform well both in the presence and the absence of PDL and thus prefer pD. Moreover, pD eliminates the problem of choosing two codes altogether making it trivial to combine immediately with any single off-the-shelf coded modulation scheme.
Fig. 1 also highlights some subtle points discussed: While iZ and D are almost indistinguishable as anticipated, D is very slightly, but measurably, better than iZ at very low SNRs since $\theta=0$ is no longer the worst angle. In particular, seeing $1+\alpha$ and $1-\alpha$ gains becomes better than seeing uniform $1-\alpha^{2}$ gains when $\theta=\pi/4$ since $1+\alpha$ acts as a diversity gain. Scheme D then has an advantage in this new worst case. While not particularly important here, this becomes relevant in other contexts. For example, if a spatially-coupled code with universality properties [18, 19] is used, theory suggests that D will be near-optimal since such codes would be agnostic to the gain imbalance as long as the mutual information is sufficient. If realized, this would eliminate the need for precoding and thus eliminate the peak-to-average power ratio (PAPR) cost of precoding.
Lastly, we note that both pD and the scheme of [1] suffer a performance loss in Fig. 1 due to error propagation in the interference cancellation. This loss is virtually eliminated if interference cancellation is done after decoding a hypothetical concatenated outer code and the performance is considered at lower BERs. This is also a practical possibility and increases the gains of pD and the scheme of [1] relative to pZ.
IV Concluding Remarks
We have illustrated that the scheme of [1] remains effective in practical coded modulation settings and provided a variation (pD) which eliminates the requirement of tuning two codes. Future work can consider alternative approaches such as exploiting the universality properties of spatially-coupled codes [18] to eliminate the need for precoding altogether.
Acknowledgments
The authors would like to acknowledge Dr. Hamid Ebrahimzad for suggesting a variation of [1] similar to pD and pointing out its practical benefits.

### v2

@@
-Polarization-Dependent Loss Mitigation with Orthogonal-Design-Based Precoding and Interference Cancellation
+Polarization-Dependent Loss Mitigation via Orthogonal Design Precoding and Interference Cancellation
 Abstract
 Recent work by Shehadeh and Kschischang provides a simple capacity-achieving scheme for channels with polarization-dependent loss (PDL) under common modeling assumptions via a careful choice of orthogonal-design-based precoding and interference cancellation.
-This letter extends that work with a simulation-based demonstration showing that this scheme remains highly effective at mitigating PDL in the highly practical setting of 4-PAM with Chase-decoded extended Hamming inner codes rather than the near-capacity inner codes considered in the original work.
-An alternative near-optimal variation of this scheme is provided requiring only one inner code rather than two and suffering no penalty in the absence of PDL, making it much more practical.
+This letter extends that work with a simulation-based demonstration showing that this scheme remains highly effective at mitigating PDL in the highly practical setting of 16-QAM with Chase-decoded extended Hamming inner codes rather than the near-capacity inner codes considered in the original work.
+An alternative near-optimal variation of this scheme is also provided requiring only one inner code rather than two and suffering no penalty in the absence of PDL, making it much more practical.
 Index Terms:
@@
 Such a model is considered in [1, 2, 3, 4, 5, 6, 7, 8, 9] and it is shown by Shehadeh and Kschischang in [1] that a certain orthogonal-design-based precoding and interference cancellation scheme is capacity-achieving under this model.
-We further assume in this letter real-valued channels with real-valued inputs.
-However, we emphasize that everything presented generalizes to complex-valued channels and inputs by analogously considering the complex variant of the scheme included in [1].
-We then have a two-parameter channel class
-$\underbrace{\begin{bmatrix}Y_{1}\\ Y_{2}\end{bmatrix}}_{\mathbf{Y}}=\underbrace{\begin{bmatrix}\sqrt{1+\gamma}&0\\ 0&\sqrt{1-\gamma}\end{bmatrix}}_{\mathbf{D}_{\gamma}}\underbrace{\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}}_{\mathbf{R}_{\theta}}\underbrace{\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}}_{\mathbf{X}}+\underbrace{\begin{bmatrix}Z_{1}\\ Z_{2}\end{bmatrix}}_{\mathbf{Z}}$
-where $\gamma\in[-\alpha,\alpha]$ and $\theta\in[0,2\pi)$, representing an adversarial class of channels with two polarizations and up to
+By a standard reduction technique, the channel matrix under this model is given by
+$\underbrace{\begin{pmatrix}\sqrt{1+\gamma}&0\\ 0&\sqrt{1-\gamma}\end{pmatrix}}_{\mathbf{D}_{\gamma}}\underbrace{\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}}_{\mathbf{R}_{\theta}}\underbrace{\begin{pmatrix}e^{i\phi}&0\\ 0&e^{-i\phi}\end{pmatrix}}_{\mathbf{B}_{\phi}}$
+which yields a three-parameter channel class
+$\mathbf{Y}=\mathbf{D}_{\gamma}\mathbf{R}_{\theta}\mathbf{B}_{\phi}\mathbf{X}+\mathbf{Z}$
+where $\gamma\in[-\alpha,\alpha]$, $\theta\in[0,2\pi)$, and $\phi\in[0,2\pi)$. This represents an adversarial class of channels corresponding to two polarizations with up to
 $10\log_{10}\left(\frac{1+\alpha}{1-\alpha}\right)\text{ dB}$
@@
 Moreover, $\mathbf{X}$ and $\mathbf{Z}$ are independent with $\mathbf{Z}$ being standard white Gaussian and $\mathbf{X}$ satisfying $\mathbb{E}[\norm{\mathbf{X}}_{2}^{2}]=2\cdot\mathsf{SNR}$ where $\mathsf{SNR}$ denotes the signal-to-noise ratio (SNR).
-We are generally interested in the worst-case performance of any given scheme over all $\gamma\in[-\alpha,\alpha]$ and $\theta\in[0,2\pi)$.
-When precoding, we consider a two-channel-use extension of the channel matrix given by $\operatorname{diag}(\mathbf{D}_{\gamma}\mathbf{R}_{\theta},\mathbf{D}_{\gamma}\mathbf{R}_{\theta})\in\mathbb{R}^{4\times 4}$ which we can interpret either as representing two uses of the channel in time or the use of a complex-valued input with real-valued representations where the first and second halves of a vector represent the real and imaginary parts respectively.
-We then take the input to the channel to be $\mathbf{X}=\mathbf{G}\mathbf{U}$ where $\mathbf{X},\mathbf{U}\in\mathbb{R}^{4\times 1}$ and $\mathbf{G}\in\mathbb{R}^{4\times 4}$ is an orthogonal precoding matrix.
-The effective channel is then described by the effective channel matrix $\operatorname{diag}(\mathbf{D}_{\gamma}\mathbf{R}_{\theta},\mathbf{D}_{\gamma}\mathbf{R}_{\theta})\mathbf{G}$ and has input $\mathbf{U}$ satisfying $\mathbb{E}[\norm{\mathbf{U}}_{2}^{2}]=2\cdot\mathsf{SNR}$.
-We use bit-interleaved coded modulation (BICM) [10] with extended Hamming codes and a Gray-labeled $4$-PAM constellation together with soft-decision Chase decoding [11].
+We are generally interested in the worst-case performance of any given scheme over all $\gamma\in[-\alpha,\alpha]$, $\theta\in[0,2\pi)$, and $\phi\in[0,2\pi)$.
+A common simplification whose justification is discussed in [2] is to take $\phi=0$ and thus remove $\mathbf{B}_{\phi}$ from the model.
+In this paper, we will cover both the simplified case of $\phi=0$ and the general case of adversarial $\phi\in[0,2\pi)$.
+In the case of $\phi=0$, it suffices to precode only across in-phase (I) and quadrature (Q) components and not across time, while in the general case of $\phi\in[0,2\pi)$, we must also precode across time.
+We will initially consider the case of $\phi=0$ for the purposes of exposition and because it is of interest in its own right, with the case of $\phi\in[0,2\pi)$ deferred to the end.
+We will work with real-valued equivalent representations of complex-valued vectors and matrices where the first and second halves of the real-equivalent of a vector contain its real and imaginary parts respectively.
+This simplifies the description of I/Q precoding and the necessary widely-linear processing to equivalent linear processing on the concatenated real–imaginary representations.
+When $\phi=0$, the channel matrix is already real-valued and its real-equivalent representation is thus $\operatorname{diag}(\mathbf{D}_{\gamma}\mathbf{R}_{\theta},\mathbf{D}_{\gamma}\mathbf{R}_{\theta})\in\mathbb{R}^{4\times 4}$.
+We take the input to the channel to be $\mathbf{X}=\mathbf{G}\mathbf{U}$ where $\mathbf{X},\mathbf{U}\in\mathbb{R}^{4\times 1}$ and $\mathbf{G}\in\mathbb{R}^{4\times 4}$ is an orthogonal precoding matrix.
+The effective channel is then described by the $4\times 4$ effective channel matrix $\operatorname{diag}(\mathbf{D}_{\gamma}\mathbf{R}_{\theta},\mathbf{D}_{\gamma}\mathbf{R}_{\theta})\mathbf{G}$ and has input $\mathbf{U}$ satisfying $\mathbb{E}[\norm{\mathbf{U}}_{2}^{2}]=4\cdot\mathsf{SNR}$.
+We use bit-interleaved coded modulation (BICM) [10] with extended Hamming codes and a $16$-QAM constellation.
+This corresponds to $\{-3,-1,1,3\}$-signaling in each degree of freedom under our real-equivalent representation.
+We further use soft-decision Chase decoding [11] in conjunction with all schemes considered.
 Since the scheme of [1] requires two codes with different rates used in equal proportion, we choose them to have the same average rate as a third code which is used in schemes requiring one code.
@@
 $\frac{1}{2}\left(\frac{k_{1}}{n_{1}}+\frac{k_{2}}{n_{2}}\right)=\frac{k}{n}\text{.}$
-We take $(n_{1},k_{1})=(100,90)$, $(n_{2},k_{2})=(400,388)$, and $(n,k)=(200,187)$.
+We take $(n_{1},k_{1})=(96,86)$, $(n_{2},k_{2})=(384,372)$, and $(n,k)=(192,179)$.
 These codes are obtained by shortening longer extended Hamming parent codes.
-This yields a rate of $0.935$ so that the inner coding scheme is somewhat comparable to that in the 400ZR implementation agreement [12].
+This yields a rate of approximately $0.93$ so that the inner coded modulation scheme altogether is somewhat comparable to that in the 400ZR implementation agreement [12].
 Our key contributions are the demonstration of the effectiveness of the scheme of [1] for PDL mitigation in this context and the introduction of a variation on this scheme with various practical advantages.
-We illustrate our points by comparing five schemes defined in Section II which include the scheme of [1] and the proposed variation.
-This is followed by simulations in Section III and concluding remarks in Section IV.
+We illustrate our points by comparing five schemes defined in Section II which include the scheme of [1] and the proposed variation followed by simulations in Section III.
+We then address more general scenarios in Section IV before concluding in Section V.
 II Five Schemes
@@
 We use zero-forcing (ZF) instead of LMMSE for simplicity since it is comparable at the high SNRs of interest.
-Implementation details for all schemes can be found in our complete simulation code which we make available online [14].
+We will assume $\phi=0$ and I/Q precoding as described in Section I for the purposes of describing our five schemes in this section.
+Implementation details for all schemes in both the $\phi=0$ case and the general $\phi\in[0,2\pi)$ case can be found in our complete simulation code which we make available online [14].
 II-A One Code, Spatio-Temporal Interleaving, and ZF without Precoding (iZ)
@@
 $\mathbf{G}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&0&1&0\\ 0&1&0&1\\ 0&1&0&-1\\ -1&0&1&0\end{pmatrix}\text{.}$ (1)
-This is followed by ZF and independent decoding of each codeword which sees a worst-case effective gain of $1-\alpha^{2}$. The resulting scheme is comparable to those of [15, 9].
+This is followed by ZF and independent decoding of each codeword which sees a worst-case effective gain of $1-\alpha^{2}$.
+The resulting scheme is essentially equivalent to those of [15, 9] as shown in [1].
+In particular, I/Q and polarization–time coding are performance-equivalent when $\phi=0$.
 II-C One Code and D-BLAST-Style ZF-SIC without Precoding (D)
 In Scheme D, we consider a D-BLAST-style [13, Chapter 8] approach which allows us to perform ZF-SIC on one codeword by diagonally staggering it in time and space.
-After mapping to a $4$-PAM constellation, we have a codeword of $n/2$ symbols which is split into two halves $\mathbf{x}^{\mathsf{f}}$ and $\mathbf{x}^{\mathsf{s}}$ of $n/4$ symbols.
+After Gray-mapping to $\{-3,-1,1,3\}$, we have a codeword of $n/2$ symbols which is split into two halves $\mathbf{x}^{\mathsf{f}}$ and $\mathbf{x}^{\mathsf{s}}$ of $n/4$ symbols.
 We then transmit $\zeta$ such codewords in $(\zeta+1)n/4$ channel uses in time which we represent by the $42\times(\zeta+1)n/$ 4 matrix
@@
 This leads to the next scheme.
-Figure 1: Simulation results for Schemes iZ, pZ, D, pD and the scheme of [1] for $\alpha=0$ and $\alpha=0.599$ corresponding respectively to worst-case PDLs of 00 dB and $6$ dB.
+Figure 1: Simulation results for Schemes iZ, pZ, D, pD, and the scheme of [1] for $\alpha=0$ and $\alpha=0.599$ corresponding respectively to worst-case PDLs of 00 dB and $6$ dB.
 II-D One Code and D-Blast-Style ZF-SIC with Precoding (pD)
@@
 In particular, we consider a two-channel-use extension and the precoding matrix (1) after which we apply Scheme D to the resulting effective channel.
-After mapping to a $4$-PAM constellation, we have a codeword of $n/2$ symbols which we split into two halves $\mathbf{u}^{\mathsf{f}}$ and $\mathbf{u}^{\mathsf{s}}$ of dimensions $82\times n/$ 8 containing $n/4$ symbols each.
+After Gray-mapping to $\{-3,-1,1,3\}$, we have a codeword of $n/2$ symbols which we split into two halves $\mathbf{u}^{\mathsf{f}}$ and $\mathbf{u}^{\mathsf{s}}$ of dimensions $82\times n/$ 8 containing $n/4$ symbols each.
 We then transmit $\zeta$ such codewords in $(\zeta+1)n/8$ uses of the two-channel-use extended channel so that the input is represented by the $84\times(\zeta+1)n/$ 8 matrix
@@
 III Simulations
+Figure 2: Simulation results for Schemes iZ, pZ, pD, and the scheme of [1] for $\alpha=0$ and $\alpha=0.599$ corresponding respectively to worst-case PDLs of 00 dB and $6$ dB and in the case of a general strictly complex-valued channel.
 We now provide bit error rate (BER) versus SNR simulation results for each of the five schemes.
@@
 This is also a practical possibility and increases the gains of pD and the scheme of [1] relative to pZ.
-IV Concluding Remarks
+IV Further Results
+When $\phi\in[0,2\pi)$, the channel matrix is complex-valued and has $4\times 4$ real-equivalent representation given by
+[…]
+which we extend to $\operatorname{diag}(\mathbf{H}_{\gamma,\theta,\phi},\mathbf{H}_{\gamma,\theta,\phi})\in\mathbb{R}^{8\times 8}$ representing two channel uses in time.
+We then take the input to be $\mathbf{X}=\mathbf{G}\mathbf{U}$ where $\mathbf{X},\mathbf{U}\in\mathbb{R}^{8\times 1}$ and $\mathbf{G}\in\mathbb{R}^{8\times 8}$ is our orthogonal precoding matrix.
+This yields an effective $8\times 8$ channel matrix given by $\operatorname{diag}(\mathbf{H}_{\gamma,\theta,\phi},\mathbf{H}_{\gamma,\theta,\phi})\mathbf{G}$ and having input $\mathbf{U}$ satisfying $\mathbb{E}[\norm{\mathbf{U}}_{2}^{2}]=8\cdot\mathsf{SNR}$.
+We then consider the $8\times 8$ precoding matrix from [1]
+$\mathbf{G}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&0&0&0&1&0&0&0\\ 0&0&1&0&0&0&1&0\\ 0&-1&0&0&0&-1&0&0\\ 0&0&0&-1&0&0&0&-1\\ 0&0&1&0&0&0&-1&0\\ -1&0&0&0&1&0&0&0\\ 0&0&0&1&0&0&0&-1\\ 0&-1&0&0&0&1&0&0\end{pmatrix}$ (2)
+and straightforwardly apply the schemes of Section II with the effective channel split into two equal halves as before.
+The complete implementation details are available in our complete simulation code [14].
+Simulation results are provided in Fig.
+2 and are nearly identical to the case of $\phi=0$ as predicted by [1].
+Scheme D is omitted since it was merely pedagogical.
+Figure 3: Simulation results for Schemes iZ, pZ, pD, and the scheme of [1] for $\alpha=0$ and $\alpha=0.333$ corresponding respectively to worst-case PDLs of 00 dB and $3$ dB and in the case of a general strictly complex-valued channel.
+Lastly, we consider the case of $3$ dB of worst-case PDL in Fig.
+3.
+In this case, we change the parameters of the code pair used in the scheme of [1] to $(n_{1},k_{1})=(144,131)$ and $(n_{2},k_{2})=(288,275)$ which have the same average rate as before but are closer to each other to account for the smaller PDL.
+Note that the slightly worse performance of the scheme of [1] in this case relative to pD is not a contradiction of the information-theoretic optimality of the scheme of [1].
+This is a finite blocklength effect due to the worse gap to capacity of the lower rate Hamming code in the code pair used.
+The scheme of [1] is always better if all inner codes used are at the same average gap to capacity.
+However, it is typically harder to achieve the same gap to capacity at lower rates that are still greater than $1/2$.
+This further highlights the significant benefits of the proposed scheme pD which obviates all of these issues and does not need to be adapted to the worst-case PDL value.
+V Concluding Remarks
 We have illustrated that the scheme of [1] remains effective in practical coded modulation settings and provided a variation (pD) which eliminates the requirement of tuning two codes.